OpenStudy (anonymous):

x + y = k x - y = k The solution to the system shown is _____. (0,k) (2, 2) (k, 0)

OpenStudy (anonymous):

Add both the equations. (x + y ) + ( x - y) = k + k 2x = 2k Therefore, x = k. Now, solve for y.

OpenStudy (anonymous):

Wait, you don't even have to solve it because of the options. -.-'

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks